Prabhleen:

The total cost for a bucket of popcorn and 4 movie tickets is $56. The total cost for the same size bucket of popcorn and 6 movie tickets is $80. The cost of a bucket of popcorn is $8. Which equation represents the relationship between y, the total cost of the popcorn and movie tickets, and x, the number of movie tickets that are purchased?

Mercury:

answer choices: y=12x+8 y=12x+24 y=4x+6 y=14x+28

Mercury:

the question isn't worded well but I believe it's asking for the price (y) of (x) movie tickets, along with one bucket of popcorn ($8). therefore, we need to find the price of one movie ticket. total price = (price of 1 popcorn bucket)(# of buckets) + (price of one ticket)(# of tickets) we are told that 1 bucket of popcorn ($8) and 4 tickets cost $56. if we let t = price of one ticket, then 56 = 8 + 4t solve for t. then, go back to the original equation, plug in the t-value for the price of tickets to get the equation in terms of x.